OpenStudy (anonymous):

1. Find the equation whose diameter has endpoints of (0, 0) and (3,3)? 2. Find the equation whose diameter has endpoints of (0,0) and (-3, 5)?

OpenStudy (anonymous):

\[\left( x-x1 \right)\left( x-x2 \right)+\left( y-y1 \right)\left( y-y2 \right)=0\]

OpenStudy (anonymous):

So I just plug the points in and that is the equation?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

yes .It is the equation of circle whose co-ordinates of diameter are given.

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (anonymous):

yw